OpenStudy (anonymous):

SOLVE 7.1=-2X-4.1(3+2.4X)

OpenStudy (anonymous):

7.1=-2X-4.1(3+2.4X) is the same has 7/10=-2x-[3+(24)/10]4/10 you know that right ?

OpenStudy (anonymous):

and then you factorize the 10's like this: 7/10=={-20x-[3+4(24)/10]}/10 \[\frac{ a }{ c }=\frac{ b }{ c }\rightarrow a=b\]

OpenStudy (anonymous):

sorry my bad 7/10=={4[-2x-[3+(24)/10]]}/10 and then 7=-8x-4[3+(24)]/10 so -8x-(12+96)/10=7 -8x=7+(12+96)10 -8x=7+(108)/10 -8x=(70+108)/10 -8x=(178)/10 x=(178)/(10*8) x=178/80 x=2.225